# Empowering 3D Visual Grounding with Reasoning Capabilities

### v2

@@ 4.3 Chain-of-Grounding Mechanism
Existing design conduct the reasoning and grounding process sequentially, i.e., the reasoning process is finished before grounding. We argue that the grounding results can also facilitate the reasoning process, especially for those requiring spatial information. Thus, to further synergize the reasoning and grounding process, inspired by chain-of-thought (CoT) [37], we propose Chain-of-Grounding (CoG), which introduces a chain of interleaved steps of reasoning and grounding to find the targeted objects during inference, as shown in Fig. 4. Such a process allows the model to actively find relevant objects that help solve the problem, and then conduct reasoning with the assistance of the additional information of these relevant objects so that the model can more precisely find the target objects.
Specifically, given a question provided by users, CoG translates it into another question of finding the explicit mentioned objects in the original question. The generated new question is sent into ReGround3D to ground the objects mentioned in the original question in the 3D scene with corresponding confidence scores. An object can be seen as a successfully located object when its confidence score is above the threshold, and the located object information could serve as explicit guidance for 3D-LLM in the next reasoning stage. As shown in Fig. 4, after obtaining the 3D information of objects explicitly mentioned in the original question, the located object information is inserted to update the question. The updated question is then sent to ReGround3D to perform reasoning and grounding to output the target object locations.
@@ 4.4.1 Training Objective
We use the pretrained weights of 3D-LLM as the initialization for the visual-centric reasoning module. Except for freezing the 3D visual encoder pretrained on [36], the rest of parameters in the visual-centric reasoning module and 3D grounding module in our framework are trained in an end-to-end manner. The training supervision is a weighted sum of the next token prediction loss from 3D-LLM and the 3D detection loss from the 3D grounding module.
@@
$\mathcal{L}_{det}=\lambda_{L1}\mathcal{L}_{L1}\\ +\lambda_{IOU}\mathcal{L}_{IOU}\\ +\lambda_{contrast}\mathcal{L}_{contrast}$ (2)
@@ 4.4.2 Instruction Tuning Dataset
We load the pretrained weights of 3D-LLM and the 3D visual encoder, and finetunes the LoRA of 3D-LLM, the query selection module, and the 3D box decoder with an instruction tuning dataset. To construct the instruction tuning dataset, we reformulate the data annotations from existing 3D datasets into question-answer or question-answer-bounding-box pairs. Specifically, the 3D visual grounding data from ScanRefer [6], SR3D, NR3D [1] and the 3D object detection data from EmbodiedScan [36] are formulated into question-answer-bbox pairs, and the 3D/spatial question answering data from SR3D [1], CLEVR3D [40], SQA3D [28] are formulated into question-answer pairs without bounding boxes. The information shown in Tab. 1 illustrates how we unify the instruction and output with task-specific templates. More details can be found in the Appendix. The reformulated data combined with our proposed ScanReason dataset, serve as the instruction tuning dataset.
@@
3D Object Detection <scene> Where is the <category> in this 3D scene? Please answer the question only
with output 3D box prediction(s). Sure, <LOC>.
@@ 5.1.1 Network Architecture
For the 3D grounding module, we adopt the pre-trained multi-modal 3D encoder in EmbodiedScan as the 3D visual encoder. During the trainning stage, We use LoRA to effeciently finetune the 3D-LLM to preserve the original 3D scene understanding capability and reduce the computation costs. The number of object queries $k$ in the query selection module is set to be 256.
@@ 5.1.2 Training Parameters
The training is done on 8 NVIDIA A100 GPUs. We adopt AdamW optimizer with a learning rate of 1e-4 and use a learning rate scheduler WarmupDecayLR with the warmup steps of 100. The total batch size is set to be 16. The loss weight parameters $\lambda_{text}$ and $\lambda_{det}$ in total loss $\mathcal{L}$ are set to 1.0 and 1.0, respectively, and the weight $\lambda_{L1}$, $\lambda_{IOU}$ and $\lambda_{contrast}$ in $\mathcal{L}_{det}$ are set to 1.0, 2.0 and 1.0.
@@ 5.2.1 Evaluation on 3D Visual Grounding
In order to verify the superiority of our model in grounding ability and facilitate comparison between current models, we report the explicit grounding performance on the existing 3D visual grounding task. Since the evaluation settings of Nr3D and Sr3D [1] are based on ground-truth object proposals, while ScanRefer [6] requires models to output 3D bounding boxes, we choose ScanRefer as our benchmark for comparison. We divide the existing methods into two categories, one is the grounding model designed specifically for the 3D visual grounding task, and the other is generalist MLLMs which can understand a variety of 3D vision-language tasks. The original 3D-LLM embed 3D locations in the vocabularies and represent the grounded 3D bounding boxes by a sequence of discrete location tokens. However, since the fine-tuned model of 3D-LLM on ScanRefer and related location tokens implementation are not accessible, we adapt 3D-LLM to directly output 3D numerical coordinates representing 3D bounding boxes by fine-tuning the pre-trained model on our reformulated 3D visual grounding data, denotes as 3D-LLM (vg). As shown in Tab. 2, current generalist MLLM models still lag behind the specialist models in terms of the grounding ability. By incorporating the 3D grounding module into MLLM, ReGround3D shows the SOTA performance on the traditional 3D visual grounding task.
@@ 5.2.2 Evaluation on 3D Reasoning Grounding
The various visual grounding models rely on explicit text-object alignment in the input object expression to achieve localization, which fall to be applied to 3D reasoning grounding task. We performed a comparision between our proposed method ReGround3D and existing MLLM methods, including 3D-LLM(vg) and Chat3D-v2 [21]. Besides, inspired by Chat-3D v2, which first segments objects, then equips them with unique object identifiers to conduct effective object grounding, we set up a LLM-based 3D reasoning grounding baseline: We first segment the objects from the scene using a 3D instance segmentor [31], then convert the segmented object information including their categories and 3D bounding boxes into text as input of LLM (InternLM2-7B [32]). Besides, to better validate the performance of our model and ensure the fair comparison, we remove the ScanReason dataset from the training data, denoted as ReGround3D*. As show in Tab. 3, the LLM-based reasoning method (Mask3D [31] + InternLM2-7B [32]) possesses a very strong function reasoning ability, but struggles understanding of 3D spatial relationship. Additionally, irrespective of whether ScanReason is used in training, our model significantly outperforms the existing MLLMs. By synergize the reasoning and grounding process utilizing the Chain-of-Grounding (CoG) mechanism, the 3D reasoning grounding performance of ReGround3D can be further improved (from 28.98 to 30.62), especially on the spatial reasoning and logistic reasoning questions. The qualitative results comparison shown in Fig. 5 demonstrates superiority of our method in reasoning human complex instruction based on the 3D scene.
@@
Methods LLM Spatial Function Logistic Emotional Safety Overall
@@ 5.3.1 Effectiveness of 3D Grounding Module
In order to more comprehensively verify the effectiveness of the 3D grounding module we proposed, we step by step verify the performance changes from 3D-LLM to ReGround3D. Apart from the 3D-LLM(vg), we fine-tuned the 3D-LLM model respectively on full reformulated existing data and all instruction tuning data, including ScanReason, denoted as 3D-LLM (full) and 3D-LLM (full+sr). All the <LOC> token in answers of the data used for fine-tuning 3D-LLM have been converted to the numerical coordinates of the corresponding boxes. The results in Tab. 4(a) show that with the same tuning dataset, our ReGround3D achieves far superior performance to 3D-LLM (28.98 vs. 19.21) by introducing the 3D grounding module.
@@ 5.3.2 Effectiveness of Instruction Tuning Dataset
Tab. 4(b) showcases the impact of different training data types on 3D visual grounding performance. 3D Object Detection (3D OD) provides the explicit 3D semantic category and visual alignment whereas 3D Question Answering (3D QA) data injects the basic 3D scene understanding ability into the model, which have certain benefits for the visual grounding ability of the model. In addition, we find that training with 3D reasoning grounding dataset can further improve the performance on 3D visual grounding.
@@ 5.3.3 Discussion of CoG Mechanism
While CoG Mechanism boosts the performance with interleaved reasoning and grounding steps during inference, it uses the relevant objects information explicitly presented in the question to help find the target objects. A natural question arises: if we input the information of all the objects instead of relavant objects into ReGround3D during CoG, will this make the model more accurately find the target objects? We first use the existing 3D Segmentor [31] to segment all objects in the scene, then update the original question using all the object 3D bounding boxes information according to the template in Sec. 4.3 and send into ReGround3D. However, experiments show that using all the object information will instead reduce the 3D reasoning grounding performance from 28.98 to 27.67. One possible reason could be the model’s attention is dispersed by too many irrelavent objects.
@@ 6 Conclusion
This paper introduces a new 3D vision language learning task: 3D reasoning grounding, which requires the model to perform active reasoning over complex and implicit human instruction, localize the target objects and give corresponding explanation. Furthermore, we propose ScanReason, a new dataset and benchmark to further unlock and throughly evaluate the 3D reasoning grounding capability. Based on this dataset, we propose a novel approach: ReGround3D, which utilizes the strong reasoning capability of MLLM guiding the 3D grounding module to obtain accurate object locations, and a Chain of Grounding (CoG) mechanism is presented to further boost the performance with interleaved reasoning and grounding steps during inference. We believe that our work will further the natural interaction between embodied agents and humans in open 3D environments. For the current ScanReason benchmark, we find that the questions in three high-level 3D reasoning categories may have overlaps. For a certain reasoning question, similar questions may appear in one or two other categories. We leave the problem as a future challenge for better reasoning grounding ability evaluation.

### v3

@@ 4.3 Chain-of-Grounding Mechanism
The existing design conducts the reasoning and grounding process sequentially, i.e., the reasoning process is finished before grounding. We argue that the grounding results can also facilitate the reasoning process, especially for those requiring spatial information. Thus, to further synergize the reasoning and grounding process, inspired by chain-of-thought (CoT) [38], we propose Chain-of-Grounding (CoG), which introduces a chain of interleaved steps of reasoning and grounding to find the targeted objects during inference, as shown in Fig. 4. Such a process allows the model to actively find relevant objects that help solve the problem, and then conduct reasoning with the assistance of the additional information of these relevant objects so that the model can more precisely find the target objects.
Specifically, given a question provided by users, CoG translates it into another question of finding the explicitly mentioned objects in the original question. The generated new question is sent into ReGround3D to ground the objects mentioned in the original question in the 3D scene with corresponding confidence scores. An object can be seen as a successfully located object when its confidence score is above the threshold, and the located object information could serve as explicit guidance for 3D-LLM in the next reasoning stage. As shown in Fig. 4, after obtaining the 3D information of objects explicitly mentioned in the original question, the located object information is inserted to update the question. The updated question is then sent to ReGround3D to perform reasoning and grounding to output the target object locations.
@@ 4.4.1 Training Objective
We use the pre-trained weights of 3D-LLM as the initialization for the visual-centric reasoning module. Except for freezing the 3D visual encoder pre-trained on [37], the rest of the parameters in the visual-centric reasoning module and 3D grounding module in our framework are trained in an end-to-end manner. The training supervision is a weighted sum of the next token prediction loss from 3D-LLM and the 3D detection loss from the 3D grounding module.
@@
$\mathcal{L}_{det}=\lambda_{IOU}\mathcal{L}_{IOU}\\ +\lambda_{contrast}\mathcal{L}_{contrast}$ (2)
@@ 4.4.2 Instruction Tuning Dataset
We load the pre-trained weights of 3D-LLM and the 3D visual encoder, and finetune the LoRA of 3D-LLM, the query selection module, and the 3D box decoder with an instruction tuning dataset. To construct the instruction tuning dataset, we reformulate the data annotations from existing 3D datasets into question-answer or question-answer-bounding-box pairs. Specifically, the 3D visual grounding data from ScanRefer [7], SR3D, NR3D [1] and the 3D object detection data from EmbodiedScan [37] are formulated into question-answer-bbox pairs, and the 3D/spatial question answering data from SR3D [1], CLEVR3D [41], SQA3D [29] are formulated into question-answer pairs without bounding boxes. The information shown in Tab. 1 illustrates how we unify the instruction and output with task-specific templates. More details can be found in the Appendix. The reformulated data combined with our proposed ScanReason dataset, serve as the instruction tuning dataset.
@@
3D Object Detection <scene> Where is the <category> in this 3D scene? Please answer the question only with output 3D box prediction(s). Sure, <LOC>.
@@ 5.1.1 Network Architecture
For the 3D grounding module, we adopt the pre-trained point cloud encoder as the 3D visual encoder. During the training stage, We use LoRA to efficiently finetune the 3D-LLM to preserve the original 3D scene understanding capability and reduce the computation costs. The number of object queries $k$ in the query selection module is set to 256.
@@ 5.1.2 Training Parameters
The training is done on 8 NVIDIA A100 GPUs. We adopt the AdamW optimizer with a learning rate of 3e-4 and use a learning rate scheduler WarmupDecayLR with the warmup steps of 100. The total batch size is set to be 16. The loss weight parameters $\lambda_{text}$ and $\lambda_{det}$ in total loss $\mathcal{L}$ are set to 1.0 and 1.0, respectively, and the weight $\lambda_{IOU}$ and $\lambda_{contrast}$ in $\mathcal{L}_{det}$ are set to 1.0 and 1.0.
@@ 5.2.1 Evaluation on 3D Visual Grounding
In order to verify the superiority of our model in grounding ability and facilitate comparison between current models, we report the explicit grounding performance on the existing 3D visual grounding task. Since the evaluation settings of Nr3D and Sr3D [1] are based on ground-truth object proposals, while ScanRefer [7] requires models to output 3D bounding boxes, we choose ScanRefer as our benchmark for comparison. We divide the existing methods into two categories, one is the grounding model designed specifically for the 3D visual grounding task, and the other is generalist MLLMs which can understand a variety of 3D vision-language tasks. The original 3D-LLM embeds 3D locations in the vocabularies and represents the grounded 3D bounding boxes by a sequence of discrete location tokens. However, since the fine-tuned model of 3D-LLM on ScanRefer and related location tokens implementation are not accessible, we adapt 3D-LLM to directly output 3D numerical coordinates representing 3D bounding boxes by fine-tuning the pre-trained model on our reformulated 3D visual grounding data, denotes as 3D-LLM (vg). As shown in Tab. 2, current generalist MLLM models still lag behind the specialist models in terms of grounding ability. By incorporating the 3D grounding module into MLLM, ReGround3D shows the SOTA performance on the traditional 3D visual grounding task.
@@ 5.2.2 Evaluation on 3D Reasoning Grounding
The various visual grounding models rely on explicit text-object alignment in the input object expression to achieve localization, which falls to be applied to 3D reasoning grounding task. We performed a comparison between our proposed method ReGround3D and existing MLLM methods, including 3D-LLM(vg) and Chat3D-v2 [22]. Besides, inspired by Chat-3D v2, which first segments objects, then equips them with unique object identifiers to conduct effective object grounding, we set up a LLM-based 3D reasoning grounding baseline: We first segment the objects from the scene using a 3D instance segmentor [32], then convert the segmented object information including their categories and 3D bounding boxes into text as input of LLM (InternLM2-7B [33]). Besides, to better validate the performance of our model and ensure a fair comparison, we remove the ScanReason dataset from the training data, denoted as ReGround3D*. As shown in Tab. 3, the LLM-based reasoning method (Mask3D [32] + InternLM2-7B [33]) possesses a very strong functional reasoning ability, but struggles to understand of 3D spatial relationship. Additionally, irrespective of whether ScanReason is used in training, our model significantly outperforms the existing MLLMs. By synergizing the reasoning and grounding process utilizing the Chain-of-Grounding (CoG) mechanism, the 3D reasoning grounding performance of ReGround3D can be further improved (from 28.98 to 30.62), especially on the spatial reasoning and logical reasoning questions. The qualitative results comparison shown in Fig. 5 demonstrates the superiority of our method in reasoning human complex instruction based on the 3D scene.
@@
Methods LLM Spatial Functional Logical Emotional Safety Overall
@@ 5.3.1 Effectiveness of 3D Grounding Module
In order to more comprehensively verify the effectiveness of the 3D grounding module we proposed, we step by step verify the performance changes from 3D-LLM to ReGround3D. Apart from the 3D-LLM(vg), we fine-tuned the 3D-LLM model respectively on full reformulated existing data and all instruction tuning data, including ScanReason, denoted as 3D-LLM (full) and 3D-LLM (full+sr). All the <LOC> tokens in answers of the data used for fine-tuning 3D-LLM have been converted to the numerical coordinates of the corresponding boxes. The results in Tab. 4(a) show that with the same tuning dataset, our ReGround3D achieves far superior performance to 3D-LLM (28.98 vs. 19.21) by introducing the 3D grounding module.
@@ 5.3.2 Effectiveness of Instruction Tuning Dataset
Tab. 4(b) showcases the impact of different training data types on 3D visual grounding performance. 3D Object Detection (3D OD) provides the explicit 3D semantic category and visual alignment whereas 3D Question Answering (3D QA) data injects the basic 3D scene understanding ability into the model, which has certain benefits for the visual grounding ability of the model. In addition, we find that training with 3D reasoning grounding dataset can further improve the performance on 3D visual grounding.
@@ 5.3.3 Discussion of CoG Mechanism
While the CoG Mechanism boosts the performance with interleaved reasoning and grounding steps during inference, it uses the relevant object information explicitly presented in the question to help find the target objects. A natural question arises: if we input the information of all the objects instead of relevant objects into ReGround3D during CoG, will this make the model more accurately find the target objects? We first use the existing 3D Segmentor [32] to segment all objects in the scene, then update the original question using all the object 3D bounding boxes information according to the template in Sec. 4.3 and send into ReGround3D. However, experiments show that using all the object information will instead reduce the 3D reasoning grounding performance from 28.98 to 27.67. One possible reason could be the model’s attention is dispersed by too many irrelevant objects.
@@ 6 Conclusion
This paper introduces a new 3D vision language learning task: 3D reasoning grounding, which requires the model to perform active reasoning over complex and implicit human instruction, localize the target objects, and give corresponding explanations. Furthermore, we propose ScanReason, a new dataset and benchmark to further unlock and thoroughly evaluate the 3D reasoning grounding capability. Based on this dataset, we propose a novel approach: ReGround3D, which utilizes the strong reasoning capability of MLLM guiding the 3D grounding module to obtain accurate object locations, and a Chain of Grounding (CoG) mechanism is presented to further boost the performance with interleaved reasoning and grounding steps during inference. We believe that our work will further the natural interaction between embodied agents and humans in open 3D environments. For the current ScanReason benchmark, we find that the questions in three high-level 3D reasoning categories may have overlaps. For a certain reasoning question, similar questions may appear in one or two other categories. We leave the problem as a future challenge for better reasoning grounding ability evaluation.
Acknowledgements
This work is supported in part by HKU Startup Fund, HKU Seed Fund for Basic Research, HKU Seed Fund for Translational and Applied Research, HKU IDS research Seed Fund, and HKU Fintech Academy R&D Funding.